OpenStudy (anonymous):

whats the reference angle for 11π/6? explain why

OpenStudy (anonymous):

The reference angle of something is its angle measure in the first quadrant. First you determine which quadrant the angle you have is in, in this case I believe it is the fourth quadrant. From there you subtract the correct angle measure to get back to the first quadrant, in this case 3pi/2 radians and that is your reference angle.

OpenStudy (campbell_st):

well this angle is in the 4th quadrant... its between \[\frac{3\pi}{2} ... and... 2\pi\] so its the same as \[2\pi - \frac{\pi}{6}\]

OpenStudy (anonymous):

thanks :) i'm not really sure what the question is asking though, is the reference angle a point or an equation?

OpenStudy (anonymous):

@azolator is it just asking the quadrant?

OpenStudy (campbell_st):

well it hase something to do with the sign of trig ratios... I'd expect.

OpenStudy (anonymous):

the reference angle is just what that angle would be if it were in the first quadrant.

OpenStudy (anonymous):

thanks :) so 3pi/2 radians is the answer? or is it more complex than that

OpenStudy (anonymous):

a little more complex than that. You need to subtract 3pi/2 from your angle so 11pi/6 - 3pi/2

OpenStudy (anonymous):

i got 1.05

OpenStudy (anonymous):

for 11pi/6 - 3pi/2 is that correct?

OpenStudy (anonymous):

@azolotor

OpenStudy (anonymous):

yeah just simplify from there

OpenStudy (anonymous):

so would that be the final answer?

OpenStudy (anonymous):

pi/3

OpenStudy (anonymous):

okay thanks :) i got 1.05 again so i guess that'd be the answer

OpenStudy (campbell_st):

here is an excellent site that explains reference angles.. http://www.mathopenref.com/reference-angle.html

OpenStudy (anonymous):

thanks :) i'll check that out. i dont really understand them too well at the moment